# Products in spinc-cobordism

### v1

@@ 3. The $Spin^{c}$ bordism ring in low degrees
Let $\tilde{T}$ denote the kernel of the ring map $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}\rightarrow\left(\Omega^{Spin^{c}}_{*}/T\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$. Filter $B$ by powers of the ideal $B\cap\tilde{T}$, i.e., equip $B$ with the $(B\cap\tilde{T})$-adic filtration. By the Anderson–Brown–Peterson splitting and by Theorem 4.2, the associated graded ring $E_{0}B$ is isomorphic to $\mathbb{F}_{2}[\beta]$ tensored with the image of $B$ in $\mathfrak{N}_{*}$ and reduced modulo the relations $\beta\cdot x=0$ for all $x\in B\cap\tilde{T}$, i.e., $E_{0}B$ is isomorphic to (9).
We claim that $B$ itself is isomorphic to (9). The $(B\cap\tilde{T})$-adic filtration on $B$ is additively split, so $B\cong E_{0}B$ as graded $\mathbb{F}_{2}$-vector spaces. In principle, the ring structure on $E_{0}B$ could differ from the ring structure on $B$ if the multiplication on $B$ were to exhibit $(B\cap\tilde{T})$-adic filtration jumps, i.e., when we multiply two elements $x,y$ of $B$, with $x$ of $B\cap\tilde{T}$-adic filtration $i$ and with $y$ of $(B\cap\tilde{T})$-adic filtration $j$, we could perhaps get an element of $(B\cap\tilde{T})$-adic filtration $>i+j$.
However, even if filtration jumps occur, $E_{0}B$ is still isomorphic to the graded $\mathbb{F}_{2}$-algebra with presentation (9). This is by a freeness argument similar to the classical argument that, if the associated graded of a filtered commutative $k$-algebra is a polynomial (i.e., free commutative) $k$-algebra, then the original filtered commutative $k$-algebra must also have been free commutative. The argument is as follows. Let $\mathcal{C}$ be the category of pairs $(A,S)$, where $A$ is a graded-commutative $\mathbb{F}_{2}[\beta]$-algebra, and $S$ is a set of homogeneous elements of $A$ such that $\beta\cdot x=0$ for all $x\in S$. There is a forgetful functor from $\mathcal{C}$ to the category $\operatorname{{\mathcal{S}\mbox{\em{ub}}}}$ of pairs $(S_{0},S_{1})$ in which $S_{0},S_{1}$ are sets and $S_{1}\subseteq S_{0}$. The forgetful functor $\mathcal{C}\rightarrow\operatorname{{\mathcal{S}\mbox{\em{ub}}}}$ sends $(A,S)$ to the underlying sets of $A$ and of $S$. The graded $\mathbb{F}_{2}[\beta]$-algebra (9) is the free object of $\mathcal{C}$ on the pair
($\left(\{y_{4},y_{8},y_{12},y_{16},\dots\}\cup\{Z_{2},Z_{6},Z_{10},Z_{18},Z_{22},\dots\},\ \{Z_{2},Z_{6},Z_{10},Z_{18},Z_{22},\dots\}\right).$
@@
In section 3, we showed that $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is not isomorphic to a polynomial algebra. Nevertheless, since an $F$-isomorphism induces a homeomorphism on the prime spectra (see [19, Proposition B.8] or [21, Lemma 29.46.9]), we have:

### v2

@@ 3. The $Spin^{c}$ bordism ring in low degrees
[…]
However, even if filtration jumps occur, $E_{0}B$ is still isomorphic to the graded $\mathbb{F}_{2}$-algebra with presentation (9). This is by a freeness argument similar to the classical argument that, if the associated graded of a filtered commutative $k$-algebra is a polynomial (i.e., free commutative) $k$-algebra, then the original filtered commutative $k$-algebra must also have been free commutative. The argument is as follows. Let $\mathcal{C}$ be the category of pairs $(A,S)$, where $A$ is a graded-commutative $\mathbb{F}_{2}[\beta]$-algebra, and $S$ is a set of homogeneous elements of $A$ such that $\beta\cdot x=0$ for all $x\in S$. There is a forgetful functor from $\mathcal{C}$ to the category $\operatorname{{\mathcal{S}\mbox{{ub}}}}$ of pairs $(S_{0},S_{1})$ in which $S_{0},S_{1}$ are sets and $S_{1}\subseteq S_{0}$. The forgetful functor $\mathcal{C}\rightarrow\operatorname{{\mathcal{S}\mbox{{ub}}}}$ sends $(A,S)$ to the underlying sets of $A$ and of $S$. The graded $\mathbb{F}_{2}[\beta]$-algebra (9) is the free object of $\mathcal{C}$ on the pair
$\left(\{y_{4},y_{8},y_{12},y_{16},\dots\}\cup\{Z_{2},Z_{6},Z_{10},Z_{18},Z_{22},\dots\},\ \{Z_{2},Z_{6},Z_{10},Z_{18},Z_{22},\dots\}\right).$
@@
In section 3, we showed that $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is not isomorphic to a polynomial algebra. Nevertheless, since an $F$-isomorphism induces a homeomorphism on the prime spectra (see [20, Proposition B.8] or [22, Lemma 29.46.9]), we have: